THE SEARCH FOR AN OBSERVABLE HELIUM COMPLEX

\underline{ADRIAN M. GARDNER}, TIMOTHY G. WRIGHT, School of Chemistry, University of Nottingham, University Park, Nottingham, NG7 2RD, United Kingdom; COREY J. EVANS, Department of Chemistry, University of Leicester, University Road, Leicester, LE1 7RH, United Kingdom.

Calculations on the He⋯MX, Ne⋯MX, and Ar⋯MX (M = Cu, Ag, Au; X = F, Cl) complexes at the CCSD and CCSD(T) levels of theory have been conducted.^{a}. The RG⋯MX (RG = He, Ne, and Ar) dissociation energies for these complexes have been evaluated by extrapolation to the complete basis set limit. The dissociation energies determined for the He⋯CuF and He⋯AuF complexes have been found to be significant, at ≈26 kJ mol^{−1}. The nature of the interactions present in these species have been investigated employing atoms-in-molecules (AIM) analysis, natural bond order analysis, and through evaluation of the dipole/induced dipole and ion/induced dipole interactions. This analysis has shown that the bonding in the strongly bound He⋯CuF and He⋯AuF complexes is slightly covalent in nature.

aC. J. Evans, T. G. Wright and A. M. Gardner, J. Phys. Chem. A, 114, 4446, (2010)